OpenStudy (anonymous):

(secA + tan A)(1-sin A)=................... Please any one to answer this question of trig identities.

OpenStudy (anonymous):

Use sec(A) + tan(A) = (1 + sin(A))/cos(A). So (sec A + tan A) (1 - sinA) = (1+sin(A))(1 -sin(A))/cos(A) = cos²(A)/cos(A) = cos(A).